1 Data Description
2 Annual Discount Rate 10%
3 Initial Investment cost – $5,000
4 Cash inflow from year 1 $1,800
5 Cash inflow from year 2 $2,100
6 Cash inflow from year 3 $1,950
7
8 Net Present Value =
Answer #1

Time n Cash Flows Present Value Factor Present Value B = (1/1.10)^n C = A*B A 0 1 2 -5,000.00 1,800.00 2,100.00 1,950.00 0.90

NPV is negative $163.04